OpenStudy (anonymous):

Which expression is the simplest form of 5/sq rt 2?

OpenStudy (danjs):

multiply by sqrt 2 over sqrt 2

OpenStudy (danjs):

\[\frac{ 5 }{ \sqrt{2} }*\frac{ \sqrt{2} }{ \sqrt{2} }\]

OpenStudy (anonymous):